OpenStudy (anonymous):

Thermometer is taken outside from a room, where the air temperature is 5F. After 1 min. the thermometer reads 55F and after 5mins it reads 30F. What is the initial temperature of the room?

OpenStudy (anonymous):

\[\frac{ dP }{ dt } = k(T-T_m)\]

OpenStudy (anonymous):

Solving the equation: \[T(t)= T_m+ce^{kt}\]

OpenStudy (anonymous):

Looks like we are looking for Tm

OpenStudy (anonymous):

If i assume T(0) = 5F outside maybe \[5 = T_m+c\] or \[5-T_m = c\]

OpenStudy (anonymous):

So at T(1) = 55 \[55 = T_m+(5-T_m)e^k\] and at T(5) = 30 \[30 = T_m+(5-T_m)e^{5k}\]

OpenStudy (anonymous):

System of equations?

OpenStudy (anonymous):

presumably, you are supposed to use Newton's law of cooling, which says that the instantaneous rate of temperature change of a body is proportional to the difference in temperature between the body and the surroundings. In mathematical form, this is expressed as: dT(t)/dt = k*(Tsurr - T(t)) where T(t) is the temperature of the body at time t, Tsurr is the temperature of the surroundings, and k is a positive constant of proportionality. I will assume that the temperature of the surroundings is constant, and that the thermometer was in thermal equilibrium with the room before being brought outside. With assumption that Tsurr is a constant, the differential equation above is a separable equation: dT/(Tsurr - T) = k dt Integrating, we get: -ln(Tsurr - T) = k*t - c where c is the constant of integration. T(t) = Tsurr + exp(c - k*t) Remember that exp(a+b) = exp(a*exp(b), so we can write the solution as: T(t) = 5F+ C*exp(-k*t) where C = exp(c), and Tsurr = 5F This is the general solution to the differential equation. It contains two unknown constants (C, and k) , but we are given two "facts" that will allow us to solve for these constants. At t = 1 min, T(1 min) = 55F and at t = 5 min, T(5min) = 30 F, so using the data at 1 min, we have: 55F = 5F+ C*exp(-k*min) 50F*exp(k*min) = C This defines C in terms of k. Now using the second condition: 30F = 5F + C*exp(-5k*min) Plug in the expression for C as a function of k: 25F = 50F*exp(k*min)*exp(-5k*min) 1/2 = exp(-4k*min) ln(1/2) = -4k*min (ln(2))/(4min) = k = 0.1733/min So: C = 50F*exp((ln(2))/4) = 59.460 F The particular solution for this case is therefore: T(t) = 5F+ (59.460F)*exp(-0.1733*t/min) At t = 0, the temperature of the thermometer, and by implication, the temperature of the room, is: T(0) = 5F + 59.460F = 64.46F